OpenStudy (anonymous):

(cos^2(theta)*(sec(theta))*(tan^2(theta))*(csc(theta))

OpenStudy (luigi0210):

This?: \[\LARGE (cos^2\theta*sec\theta)(tan^2\theta*csc\theta)\]

OpenStudy (anonymous):